OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form

jimthompson5910 (jim_thompson5910):

y = mx+b y = -1x+ b ... plug in m = -1 y = -x + b 4 = -1*(-1)+b ... plug in x = -1 and y = 4 Now solve for b

jimthompson5910 (jim_thompson5910):

Remember to convert to standard form

OpenStudy (anonymous):

x+y=3.. wait that isn't standard form

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

that's standard form and it's 100% correct, nice work

OpenStudy (anonymous):

thanks haha I was slightly unsure of myself

jimthompson5910 (jim_thompson5910):

lol well you'll get more confident as you do these